# Preventing Model Collapse Under Overparametrization: Optimal Mixing Ratios for Interpolation Learning and Ridge Regression

### v1

@@ 3.1 Min-$\ell_{2}$-norm interpolator
Note Theorem 3.1 applies for any $\bm{\Sigma}$} obeying our assumptions. For the special case of isotropic features, i.e., ${\bm{\Sigma}}=I$, the limiting risk simplifies to
@@
Effect of number of iteration $t$ and mixing parameter $w$: In equation (3.4), the first term corresponds to the variance while the second to the bias (recall definitions from equation (2.7)). Note that the quantities $\sigma^{2}\mathcal{V}$ and $b_{\star}\mathcal{B}$ coincides asymptotic variance and bias of min-$\ell_{2}$-norm interpolators in overparametrized regression (Hastie et al., 2022, Theorem 2). In fact, the bias term $B(\widehat{\boldsymbol{\beta}}_{t},\boldsymbol{\beta})$ is independent of both the iteration $t$ and mixing proportion $w$. The impact of mixing on the generalization error arises through only the variance, captured by the function $c(w)$, which is minimized at $w^{\star}=\varphi^{-1}\approx 0.618$. This golden-ratio weighting phenomenon was previously observed for ordinary least squares in low-dimensional regression with Gaussian covariates He et al. (2025).
@@
It follows immediately that $w^{\star}_{t}$ is decreasing, so if one is free to adjust the mixing proportion at every generation, the optimal strategy places progressively more weight on the synthetic data. Moreover, using $w^{\star}_{t}\rightarrow w^{\star}$ (as defined in Theorem 3.1), in the long run, the asymptotic risk is the same whether one uses a fixed $w^{\star}$ across all generations or adapts $w^{\star}_{t}$ dynamically, as proven in Section C.
Next, we study ridge regression. Here, both the variance and bias terms will depend on $\lambda,t,w$. For the sake of clarity, we present our results for isotropic covariates in Subsection 3.2. We discuss the case of non-isotropic covariates later in Subsection $3.$
3.2 Ridge regression: Isotropic covariance (${\bm{\Sigma}}=\alpha I$)
In case of isotropic features, i.e., ${\bm{\Sigma}}=\alpha I$, the bias and variance of $\widehat{\boldsymbol{\beta}}_{t,\lambda}$ simplifies since $s_{k}\equiv\alpha$. Hence, by equation (3.1), $G=H=\delta_{\alpha}$, where $\delta_{x}$ denote the Dirac probability measure at $x$. This implies $m(z)$, from  equation (3.2), simplifies to be the unique solution to
@@ 3.2 Ridge regression: Isotropic covariance (${\bm{\Sigma}}=\alpha I$)
Assume ${\bm{\Sigma}}=\alpha I$ and the setting of Section 2. For $0<w<1$, $\lambda>0$, we have almost surely over the randomness in the covariates
$\lim_{n\to\infty}\lim_{t\to\infty}R(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})=\sigma^{2}\,c(w)\,\mathcal{V}_{\lambda}\;+\;b_{\star}\,\mathcal{B}_{\lambda},$
@@
Under the assumptions of Theorem 3.2, let $w^{\star}(\lambda)$ be the unique global minimizer of the limiting risk as defined in Theorem 3.2. Then $w^{\star}(\lambda)$ is a continuous function of $\lambda$ satisfying
  1. (i)
    $w^{\star}(\lambda)\in[0.5,1]$,
  2. (ii)
    $w^{\star}(\lambda)\to\phi^{-1}$ as $\lambda\downarrow 0$,
  3. (iii)
    $w^{\star}(\lambda)\to 1$ as $\lambda\uparrow\infty$.
Furthermore, $w^{\star}(\lambda)$ is an increasing function of SNR.
@@ 3.3 Ridge regression: Correlated features
In the setting of Section 2, suppose $w\in(0,1)$, and $\lambda>0$. Define $m_{1}=m(-\lambda/w)$, $m_{2}=m(-\lambda/(2-w))$, where $m(\cdot)$ is the unique solution to equation (3.2). Then almost surely over the randomness in the covariates,
$\lim_{n\to\infty}\lim_{t\to\infty}R(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})=\sigma^{2}\,c(w)\,\mathcal{V}_{\lambda}+b_{\star}\,\mathcal{B}_{\lambda},\quad\text{with}$ (3.8)
@@
Proof of Theorem 3.4 is available in Appendix A.
Theorem 3.4 is our most general result. It shows for any $w\in(0,1)$ and any $\lambda>0$, $R(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})$ does not diverge even when $t$ increases. The result highlights the necessity of mixing real-data with synthetic outputs to mitigate model collapse. In its most general form the risk in equation (3.8) is involved to analyze. In the following subsections, we study two popular models where the risk simplifies and the optimal mixing ratio can be studied analytically.
@@ 3.4.1 Random Effects Model
Modern applications, ranging from text classification to economics and genomics, are characterized by dense but weak signals spread across many coordinates (Joachims, 1998; Boyle et al., 2017; Yang et al., 2010; Shen & Xiu, 2025). This setting is well captured by a random-effects model, where each feature contributes a small, independent effect, and it provides a simple yet sophisticated framework for rigorously analyzing interesting high-dimensional predictors (Dobriban & Wager, 2018). Adopting a random-effects framework, we assume that each coordinate $\beta_{i}$ of the signal is drawn i.i.d. with $\mathbb{E}\beta_{i}=0$ and $\mathrm{Var}(\beta_{i})=b_{\star}/p>0$. In this setting, the limiting risk can be characterized as follows.
@@
Suppose $\beta_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}(0,b_{\star}/p)$. Assume the framework of Section 2 and fix $0<w<1$, $\lambda>0$. Let $m(\cdot)$ be the solution to equation (3.2) and define $f(z)=m(-z)^{-1}-z$. Then almost surely over the randomness in the covariates
$\lim_{n\to\infty}\lim_{t\to\infty}R(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})=\sigma^{2}\,c(w)\,\mathcal{V}_{\lambda}+b_{\star}\,\mathcal{B}_{\lambda},\quad\text{where}$
@@
The representation of risk via the risk function $f$ enables to rigorously study properties of the optimal mixing ratio, as presented below.
@@
While Propositions 3.1 and 3.2 are stated under the random effects assumption, the conclusions continue to hold for a broader class of parameters $(\boldsymbol{\beta},{\bm{\Sigma}})$. If the probability measures $\widehat{G}_{p}$ and $\widehat{H}_{p}$ from equation (3.1) converge weakly to the same distribution, i.e., $G=H$, we have the same conclusions. In fact, we prove in Lemma D.1 that the random-effects assumption can be seen as a special case of $G=H$ . Other natural examples where $G=H$ include the isotropic covariance case, ${\bm{\Sigma}}=\alpha I$, or where $\boldsymbol{\beta}$ is drawn uniformly at random from a $p$-dimensional sphere, or $\bm{\Sigma}$} is drawn from an orthogonally invariant ensemble.
@@ 3.4.2 Spiked covariance model
For our second application, we consider a popular class of covariance matrices – the spiked covariance model Birnbaum et al. (2013); Johnstone (2001); Johnstone & Onatski (2020). In the past two decades, this covariance class has seen exciting applications in population genetics Patterson et al. (2006); Price et al. (2006), finance Knight et al. (2005); Ledoit & Wolf (2022), and signal processing Johnstone & Lu (2009); Wang et al. (2024), among others.
Formally, we assume ${\bm{\Sigma}}=I+svv^{\top}$ for some $v\in\mathbb{R}^{p}$, with $\|v\|_{2}=1$, and $s>0$. The results below extend naturally to spiked models with multiple but finitely many spikes, but for simplicity, we study the problem for the case of a single spike. The limiting risk takes the following form.
@@
In the setting of Section 2, assume that ${\bm{\Sigma}}=I+svv^{\top}$ and the signal takes the form $\boldsymbol{\beta}=\theta v+\sqrt{1-\theta^{2}}v^{\perp}$, while satisfying equation (2.2), with $v^{\top}v^{\perp}=0$ and $\|v^{\perp}\|_{2}=1$. If $\theta\equiv\theta(n)\rightarrow\theta_{\star}$, then we have almost surely over the randomness of covariates that
$\lim_{n\to\infty}\lim_{t\to\infty}R(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})=\sigma^{2}\,c(w)\,\mathcal{V}_{\lambda}+b_{\star}\,\mathcal{B}_{\lambda},$
where $\mathcal{V}_{\lambda}$ is the same as in Theorem 3.2 and
@@ 4 Simulations
In this section, we conduct numerical experiments to complement our theoretical findings. For all empirical risk plots, we generate an $n\times p$ feature matrix $\mathbf{X}=\mathbf{Z}{\bm{\Sigma}}^{1/2}$, where $Z_{ij}\sim\mathcal{N}(0,1)$ and vary $\bm{\Sigma}$} across different plots. We display both the theoretical risk predicted by our formulae (solid lines) and corresponding empirical estimates ($\times$ markers). The empirical risks are calculated by averaging over $100$ runs.
@@
Risk of min-$\ell_{2}$-norm Interpolator In Figure 1, we plot the asymptotic generalization error of the min-$\ell_{2}$-norm interpolator as a function of mixing weight $w$ for two different classes of covariance matrices: i. ${\bm{\Sigma}}=I$ (Panel $(a)$) and ii. ${\bm{\Sigma}}_{ij}=2^{-|i-j|}$ (Panel $(b)$), corresponding to the correlation matrix of an AR(1)(1) process. We vary $\gamma=1.5,2,3$. The choice of the remaining parameters are as follows: sample size $n=200$, number of features $p=\gamma n$, and number of iterations $t=5$. To generate $\boldsymbol{\beta}$, we first simulate $\widetilde{\boldsymbol{\beta}}$ with $\widetilde{\beta}_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Bern}(0.1)$. Set $\boldsymbol{\beta}=\widetilde{\boldsymbol{\beta}}/\|\widetilde{\boldsymbol{\beta}}\|_{2}$ which yields $b_{\star}=1$.
@@
We observe that the empirical risk matches with its theoretical counterpart even for moderate sample size. Further, the generalization error is always minimized at $w^{\star}=\varphi^{-1}\approx 0.618$ (dashed vertical line), consistent with Theorem 3.1. Panel $(c)$ shows the risk of the min-$\ell_{2}$-norm interpolator as a function of iteration $t$. We have computed the risk at optimal mixing weight $w=\varphi^{-1}$ and ${\bm{\Sigma}}=I$. We observe that both theoretical and empirical risks stabilize after only a few iterations.
@@
Optimal weights In Figure 2, we consider several properties of generalization error of ridge estimator $\widehat{\boldsymbol{\beta}}_{t,\lambda}$. In Figure 2(a)(a), $(b)$, we plot optimal weight $w^{\star}$ as a function of $\lambda$. We set $n=200,p=\gamma n$, and $t=5$, and vary $\gamma=1.2,2,4$. Figure 2 $(a)$ considers isotropic covariance ${\bm{\Sigma}}=I$ with high noise variance $\sigma^{2}=64$. The plot demonstrates that for low SNR, $w^{\star}$ can be arbitrarily close to $0.5$. Further, $w^{\star}(\lambda)$ is neither monotone, nor convex as a function of $\lambda$ with $w^{\star}(\lambda)\rightarrow\varphi^{-1}$ as $\lambda\rightarrow 0+$.
@@
Figure 2 $(b)$ corresponds to spiked covariance model ${\bm{\Sigma}}=I+5e_{1}e_{1}^{\top}$, where $e_{1},\ldots,e_{p}$ denote the canonical basis of $\mathbb{R}^{p}$. We set $\beta_{1}=0.5$ and $\boldsymbol{\beta}_{2:p}=\sqrt{1-0.5^{2}}\times\widetilde{\boldsymbol{\beta}}/\|\widetilde{\boldsymbol{\beta}}\|$ where $\widetilde{\beta}_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Bern}(0.25)$. This implies that $\|\boldsymbol{\beta}\|^{2}=1$ and $\theta_{\star}=0.5$, where $\theta_{\star}$ defined as in Proposition 3.3. The plot shows that for large regularization parameter $\lambda$, we have $w^{\star}=1$, consistent with Proposition 3.3.
Figure 2 $(c)$ plots the risk of $\widehat{\boldsymbol{\beta}}_{t,\lambda}$ with $\sigma^{2}=1$, $w=\varphi^{-1}$. We set $\bm{\Sigma}$} to be equicorrelated, i.e.,
@@
${\bm{\Sigma}}=\left(1-\frac{\rho}{\sqrt{p}}\right)I+\frac{\rho}{\sqrt{p}}\,\mathbf{1}\mathbf{1}^{\top},\qquad\text{with }\rho=1/2,$
@@
and $\beta_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,p^{-1})$. The empirical risk is computed with $\sigma^{2}=1,t=10$, $n=400$. Note that $\bm{\Sigma}$} does not satisfy bounded eigenvalue condition in Assumption 2.1, as its largest eigenvalue $\geq C\sqrt{p}$ for some $C>0$. Nevertheless, the generalization error is accurately predicted by Proposition 3.1 as long as $\boldsymbol{\beta}$ lies in a generic position relative to $\bm{\Sigma}$}. This demonstrates the robustness of our theoretical findings.
@@
Figure 1: Generalization error of min-$\ell_{2}$-norm interpolator as a function of weight $w$ (Panel $(a)$ and $(b)$) and iterations $t$ (Panel $(c)$) for different values of $\gamma$. Panel $(a)$ considers isotropic covariance ${\bm{\Sigma}}=I$ and panel $(b)$ considers anisotropic $\bm{\Sigma}$} with ${\bm{\Sigma}}_{ij}=2^{-|i-j|}$, which corresponds to covariance matrix of AR(1)(1) model. In panes $(a)$ and $(b)$, the risk is minimized at $w^{\star}=1/\varphi$, as proved by Theorem 3.1. Panel $(c)$ shows that both empirical and theoretical risks stabilize in a few iterations.

### v2

@@ 3.1 Min-$\ell_{2}$-norm interpolator
Note Theorem 3.1 applies for any $\bm{\Sigma}$} obeying our assumptions. For the special case of isotropic features, i.e., ${\bm{\Sigma}}=\bm{I}$, the limiting risk simplifies to
@@
Effect of number of iteration $t$ and mixing parameter $w$: In equation (3.4), the first term corresponds to the variance while the second to the bias (recall definitions from equation (2.5)). Note that the quantities $\sigma^{2}\mathcal{V}$ and $b_{\star}\mathcal{B}$ coincides asymptotic variance and bias of min-$\ell_{2}$-norm interpolators in overparametrized regression (Hastie et al., 2022, Theorem 2). In fact, the bias term $B(\widehat{\boldsymbol{\beta}}_{t},\boldsymbol{\beta})$ is independent of both the iteration $t$ and mixing proportion $w$. The impact of mixing on the generalization error arises through only the variance, captured by the function $c(w)$, which is minimized at $w^{\star}=\varphi^{-1}\approx 0.618$. This golden-ratio weighting phenomenon was previously observed for a host of models in low-dimensional settings He et al. (2025).
@@
It follows immediately that $w^{\star}_{t}$ is decreasing, so if one is free to adjust the mixing proportion at every generation, the optimal strategy places progressively more weight on the synthetic data. Moreover, using $w^{\star}_{t}\rightarrow w^{\star}$ (as defined in Theorem 3.1), in the long run, the asymptotic risk is the same whether one uses a fixed $w^{\star}$ across all generations or adapts $w^{\star}_{t}$ dynamically, as proven in Appendix C. Our findings can also be seen empirically in Figure 3 (b), where we plot the risk vs number of iterations for different values of the weight $w$. The difference between $w=\varphi$ and dynamic $w$ is minimal and only visible at $t=1$. Further, this difference is minimal compared to the difference between $w=1$ and $w=\varphi^{-1}$. The empirical risks were calculated at $n=200,p=400$ with $\texttt{SNR}=1$.
Next, we study ridge regression. Here, both the variance and bias terms will depend on $\lambda,t,w$.
3.2 Ridge regression: Isotropic covariance (${\bm{\Sigma}}=\alpha\bm{I}$)
In case of isotropic features, i.e., ${\bm{\Sigma}}=\alpha\bm{I}$, the bias and variance of $\widehat{\boldsymbol{\beta}}_{t,\lambda}$ simplifies since $s_{k}\equiv\alpha$. Hence, by equation (3.1), $G=H=\delta_{\alpha}$, where $\delta_{x}$ denote the Dirac probability measure at $x$. This implies $m(z)$, from  equation (3.2), simplifies to be the unique solution to
@@ 3.2 Ridge regression: Isotropic covariance (${\bm{\Sigma}}=\alpha\bm{I}$)
Assume ${\bm{\Sigma}}=\alpha\bm{I}$ and the setting of Section 2. For $0<w<1$, $\lambda>0$, we have almost surely over the randomness in the covariates
$\lim_{n\to\infty}\lim_{t\to\infty}R(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})=\sigma^{2}\,c(w)\,\mathcal{V}_{\lambda}\;+\;b_{\star}\,\mathcal{B}_{\lambda},$ (3.7)
@@
Under the assumptions of Theorem 3.2, let $w^{\star}(\lambda)$ be the unique global minimizer of the limiting risk as defined in Theorem 3.2. Then $w^{\star}(\lambda)$ is an increasing function of SNR, a continuous function of $\lambda$ and satisfies (i) $w^{\star}(\lambda)\in[0.5,1]$, (ii) $w^{\star}(\lambda)\to\varphi^{-1}\approx 0.618$ as $\lambda\downarrow 0$ and (iii) $w^{\star}(\lambda)\to 1$ as $\lambda\uparrow\infty$.
@@ 3.3 Ridge regression: Correlated features
In the setting of Section 2, suppose $w\in(0,1)$, and $\lambda>0$. Define $m_{1}=m(-\lambda/w)$, $m_{2}=m(-\lambda/(2-w))$, where $m(\cdot)$ is the unique solution to equation (3.2). Then almost surely over the randomness in the covariates, equation (3.7) holds with
@@
The proof is deferred to Appendix A. Theorem 3.4 is our most general result. It shows for any $w\in(0,1)$ and any $\lambda>0$, $R(\widehat{\boldsymbol{\beta}}_{t,\lambda};\boldsymbol{\beta})$ does not diverge even when $t$ increases. The result highlights the necessity of mixing real-data with synthetic outputs to mitigate model collapse. In its most general form the risk in equation (3.7) is involved to analyze. In the sequel, we study two popular models where the risk simplifies and the optimal mixing ratio can be studied analytically.
@@ 3.4.1 Random Effects Model
Modern applications, ranging from text classification to economics and genomics, are characterized by dense but weak signals spread across many coordinates (Joachims, 1998; Boyle et al., 2017; Yang et al., 2010; Shen and Xiu, 2025). This setting is well captured by a random-effects model, where each feature contributes a small, independent effect, and it provides a simple yet sophisticated framework for rigorously analyzing interesting high-dimensional predictors (Dobriban and Wager, 2018). Adopting a random-effects framework, we assume that each coordinate $\beta_{i}$ of the signal is drawn i.i.d. with $\mathbb{E}\beta_{i}=0$ and $\mathrm{Var}(\beta_{i})=b_{\star}/p>0$.
@@
Suppose $\beta_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}(0,b_{\star}/p)$. Assume the framework of Section 2 and fix $0<w<1$, $\lambda>0$. Let $m(\cdot)$ be the solution to equation (3.2) and define $f(z)=m(-z)^{-1}-z$. Then almost surely over the randomness in the covariates, equation (3.7) holds with
@@
With this characterization for the risk, we show the optimal mixing ratio satisfies the following.
@@
While Propositions 3.1 and 3.2 are stated under the random effects assumption, the conclusions continue to hold for a broader class of parameters $(\boldsymbol{\beta},{\bm{\Sigma}})$. If the probability measures $\widehat{G}_{p}$ and $\widehat{H}_{p}$ from equation (3.1) converge weakly to the same distribution, i.e., $G=H$, we have the same conclusions. In fact, we prove in Lemma D.1 that the random-effects assumption can be seen as a special case of $G=H$ . Other natural examples where $G=H$ include the isotropic covariance case, ${\bm{\Sigma}}=\alpha\bm{I}$, or where $\boldsymbol{\beta}$ is drawn uniformly at random from a $p$-dimensional sphere, or $\bm{\Sigma}$} is drawn from an orthogonally invariant ensemble.
@@ 3.4.2 Spiked covariance model
For our second application, we consider a popular class of covariance matrices – the spiked covariance model Birnbaum et al. (2013); Johnstone (2001); Johnstone and Onatski (2020). In the past two decades, this covariance class has seen exciting applications in population genetics Patterson et al. (2006); Price et al. (2006), finance Knight et al. (2005); Ledoit and Wolf (2022), and signal processing Johnstone and Lu (2009); Wang et al. (2024), among others.
Formally, we assume ${\bm{\Sigma}}=\bm{I}+s\bm{v}\bm{v}^{\top}$ for some $v\in\mathbb{R}^{p}$, with $\|\bm{v}\|_{2}=1$, and $s>0$. The results below extend naturally to spiked models with multiple but finitely many spikes, but for simplicity, we study the problem for the case of a single spike. The limiting risk takes the following form.
@@
In the setting of Section 2, assume that ${\bm{\Sigma}}=\bm{I}+s\bm{v}\bm{v}^{\top}$ and the signal takes the form $\boldsymbol{\beta}=\theta\bm{v}+\sqrt{1-\theta^{2}}\bm{v}^{\perp}$, while satisfying equation (2.2), with $\bm{v}^{\top}\bm{v}^{\perp}=0$ and $\|\bm{v}^{\perp}\|_{2}=1$. If $\theta\equiv\theta(n)\rightarrow\theta_{\star}$, then we have almost surely over the randomness of covariates that equation (3.7) holds with $\mathcal{V}_{\lambda}$ same as in Theorem 3.2 and
@@ 4 Simulations
We conduct numerical experiments to complement our theoretical findings. For all plots, we generate $\mathbf{X}=\mathbf{Z}{\bm{\Sigma}}^{1/2}$, where $Z_{ij}\sim\mathcal{N}(0,1)$ and vary $\bm{\Sigma}$} across different plots. We display both the theoretical risk predicted by our formulae (solid lines) and corresponding empirical estimates ($\times$ markers). For the first two plots, the empirical risks are calculated by averaging over $100$ runs.
@@
Risk of min-$\ell_{2}$-norm Interpolator In Figure 1, we plot the asymptotic generalization error of the min-$\ell_{2}$-norm interpolator as a function of mixing weight $w$ for two different classes of covariance matrices: i. ${\bm{\Sigma}}=\bm{I}$ (Panel $(a)$) and ii. ${\bm{\Sigma}}_{ij}=2^{-|i-j|}$ (Panel $(b)$), corresponding to the correlation matrix of an AR(1)(1) process. We vary $\gamma=1.5,2,3$. The choice of the remaining parameters are as follows: sample size $n=200$, number of features $p=\gamma n$, and number of iterations $t=5$. To generate $\boldsymbol{\beta}$, we first simulate $\widetilde{\boldsymbol{\beta}}$ with $\widetilde{\beta}_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Bern}(0.1)$. Set $\boldsymbol{\beta}=\widetilde{\boldsymbol{\beta}}/\|\widetilde{\boldsymbol{\beta}}\|_{2}$ which yields $b_{\star}=1$.
@@
We observe that the empirical risk matches with its theoretical counterpart even for moderate sample size. Further, the generalization error is always minimized at $w^{\star}=\varphi^{-1}\approx 0.618$ (dashed vertical line), consistent with Theorem 3.1. Panel $(c)$ shows the risk of the min-$\ell_{2}$-norm interpolator as a function of iteration $t$. We have computed the risk at optimal mixing weight $w=\varphi^{-1}$ and ${\bm{\Sigma}}=\bm{I}$. We observe that both theoretical and empirical risks stabilize after only a few iterations.
@@
Optimal weights in ridge regression In Figure 2, we plot the optimal weight $w^{\star}$ as a function of $\lambda$. We set $n=200,p=\gamma n$, and $t=5$, and vary $\gamma=1.2,2,4$. Figure 2 $(a)$ considers isotropic covariance ${\bm{\Sigma}}=\bm{I}$ with high noise variance $\sigma^{2}=64$. The plot demonstrates that for low SNR, $w^{\star}$ can be arbitrarily close to $0.5$. Further, $w^{\star}(\lambda)$ is neither monotone, nor convex as a function of $\lambda$ with $w^{\star}(\lambda)\rightarrow\varphi^{-1}$ as $\lambda\rightarrow 0^{+}$.
@@
Figure 2 $(b)$ corresponds to the spiked covariance model ${\bm{\Sigma}}=\bm{I}+5\bm{e}_{1}\bm{e}_{1}^{\top}$, where $\bm{e}_{1},\ldots,\bm{e}_{p}$ denote canonical basis vectors. We set $\beta_{1}=0.5$ and $\boldsymbol{\beta}_{2:p}=\sqrt{1-0.5^{2}}\times\widetilde{\boldsymbol{\beta}}/\|\widetilde{\boldsymbol{\beta}}\|$ where $\widetilde{\beta}_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Bern}(0.25)$. This implies that $\|\boldsymbol{\beta}\|^{2}=1$ and $\theta_{\star}=0.5$, where $\theta_{\star}$ defined as in Proposition 3.3. The plot shows that for large regularization parameter $\lambda$, we have $w^{\star}=1$, consistent with Proposition 3.3. Figure 2 $(c)$ plots the risk of $\widehat{\boldsymbol{\beta}}_{t,\lambda}$ with $\sigma^{2}=1$, $w=\varphi^{-1}$. We set $\bm{\Sigma}$} to be equicorrelated, i.e.,
@@
${\bm{\Sigma}}=\left(1-\frac{\rho}{\sqrt{p}}\right)\bm{I}+\frac{\rho}{\sqrt{p}}\,\mathbf{1}\mathbf{1}^{\top},\qquad\text{with }\rho=1/2,$
@@
and $\beta_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,p^{-1})$. The empirical risk is computed with $\sigma^{2}=1,t=10$, $n=400$. Note that $\bm{\Sigma}$} does not satisfy bounded eigenvalue condition Assumption 2.1, as its largest eigenvalue $\geq C\sqrt{p}$ for some $C>0$. Still, the generalization error is accurately predicted by Proposition 3.1 as long as $\boldsymbol{\beta}$ lies in a generic position relative to $\bm{\Sigma}$}. This demonstrates the robustness of our theoretical findings. Figure 3 (a) plots theoretical values of $w^{\star}$ as a function of $\gamma$ for different values of $\lambda$ (we fixed $\texttt{SNR}=1/4$). As expected, when $\lambda$ is close to 0, $w^{\star}\approx\varphi^{-1}$. However, as $\lambda$ increases, the dependence on $\gamma$ becomes more visible. This dependence of the optimal weight on $\gamma$ is important to note, since it illustrates the importance of studying our high-dimensional framework.
@@
Figure 1: Generalization error of min-$\ell_{2}$-norm interpolator as a function of weight $w$ (Panel $(a)$ and $(b)$) and iterations $t$ (Panel $(c)$) for different values of $\gamma$. Panel $(a)$ considers isotropic covariance ${\bm{\Sigma}}=\bm{I}$ and panel $(b)$ considers anisotropic $\bm{\Sigma}$} with ${\bm{\Sigma}}_{ij}=2^{-|i-j|}$, which corresponds to covariance matrix of AR(1)(1) model. In panes $(a)$ and $(b)$, the risk is minimized at $w^{\star}=1/\varphi$, as proved by Theorem 3.1. Panel $(c)$ shows that both empirical and theoretical risks stabilize in a few iterations.